# Hybrid-pair superfluidity in a strongly driven Fermi gas

### v1

@@ IV.4 Thouless criterion at finite temperature
×$\times\begin{cases}\exp(\frac{\pi/c^{2}}{2k_{F}a_{13}}+\frac{\pi^{2}}{m_{r}k_{F}}\frac{s^{2}}{c^{2}}I_{+}),&a_{23}=0,\\ \exp(\frac{\pi/s^{2}}{2k_{F}a_{23}}+\frac{\pi^{2}}{m_{r}k_{F}}\frac{c^{2}}{s^{2}}I_{+}),&a_{13}=0.\end{cases}$
@@ V.1.2 Magnetization
An interesting property of the present Rabi driven Fermi superfluid is its magnetization in terms of the Rabi coupled species |$\ket{1}$ and |$\ket{2}$. The components of the magnetization vector are given by
@@ V.1.3 Two- and three-body contacts
$\leavevmode\nobreak\ \leavevmode\nobreak\ +\sum_{{\bf k}{\bf p}{\bf q}}v_{\bf k}^{2}v_{{\bf p}}u_{\bf p}v_{{\bf q}}u_{\bf q}\left(c_{\bf k}^{2}s_{\bf p}s_{\bf q}+s_{\bf k}^{2}c_{\bf p}c_{\bf q}-2s_{\bf k}c_{\bf k}s_{\bf p}c_{\bf q}\right),$
@@
Figure 6: BEC-BCS crossover with Rabi-coupled fermions. (a) Effective scattering length, (b) BCS order parameter, (c) total chemical potential measured from the (Rabi-shifted) continuum. The results are plotted for parameters $11/k_{F}a_{13}=$ 1, $a_{23}=0$, and $\Omega/E_{F}=1,2,3$ (solid, dashed, and dotted lines). The inset in (b) shows the analytical order parameter in the BCS regime, Eq. (73).
@@
where we have used the facts that $g_{j3}\rightarrow 0^{-}$ and $g_{13}/g_{23}\rightarrow 1$ when we take $\Lambda\rightarrow\infty$ in Eq. (3). Since our ansatz only explicitly involves two-body correlations, the three-body contact is entirely determined from the two-body contacts and the densities. In this sense, the three-body contact as defined above can be considered to be “disconnected” and distinct from the thermodynamic variable introduced in Refs. [64, 65], which depends on genuine three-body correlations and on physics beyond the two-body scattering length (e.g., on the existence of a three-body parameter [66]). However, the three-body contact in Eq. (70) still has an important physical interpretation as it determines the rate of three-body losses, i.e., the loss rate is directly proportional to $\mathcal{C}^{(3)}_{123}$ (see, e.g., Ref. [67]).
@@ V.2 BEC-BCS crossover for a single interaction channel
We first wish to investigate the influence of the detuning on the system properties. To achieve this, we examine the effective scattering length $a_{-}$, which characterizes interactions between particles of species |$\ket{3}$ and the lower-energy Rabi-driven superposition state of species |$\ket{1}$ and |$\ket{2}$ [41]. When $a_{23}=0$, the dressed scattering length (54) reduces to
@@
We see that when $\delta/\Omega\gg 1$, $a_{-}$ approaches the bare scattering length $a_{13}$, and that if $a_{13}>0$, $a_{-}$ can exhibit a resonance (i.e., a zero crossing of $1/a_{-}$) by controlling the detuning. This behavior is illustrated in Fig. 6(a) which shows $1/k_{F}a_{-}$ as a function of detuning for a fixed bare scattering length $11/k_{F}a_{13}=$ 1 and different magnitudes of the Rabi drive: $\Omega/E_{F}=1$, 2, and 3.
@@
Figure 7: (a) Out-of-plane and in-plane magnetization $\mathcal{S}_{z}$ (red) and $\mathcal{S}_{x}$ (blue), respectively, and (b) disconnected three-body contact. Parameters as in Fig. 6: $11/k_{F}a_{13}=$ 1, $a_{23}=0$, and $\Omega/E_{F}=1,2,3$ (solid, dashed, and dotted lines).
@@
Figure 8: (a,b,c) Density distributions and (d,e,f) pairing correlations across the BEC-BCS crossover. From left to right we have $\delta/E_{F}=-3.4,-2.1,3$. The dashed black lines in (d,e,f) are obtained using the low-density approximation for $c_{\bf k}$ and $s_{\bf k}$ when $\bar{\xi}_{\bf k}\geq|\epsilon_{-}|$ [Eq. (23)] and the noninteracting coefficients [Eq. (6)] when $\bar{\xi}_{\bf k}<|\epsilon_{-}|$, substituted into Eq. (57). The results are plotted for parameters $\Omega/E_{F}=1$, $11/k_{F}a_{13}=$ 1, $a_{23}=0$.
We also investigate the multi-component nature of the Rabi-coupled Fermi gas, as encoded in the magnetization and three-body contact. In Fig. 7(a) we see that $\mathcal{S}_{z}$ evolves smoothly from $-1$ to $+1$ as a function of detuning. This is natural as the ground state changes from being composed of a set of weakly perturbed Fermi seas in the $|2\rangle$ and $|3\rangle$ components at large negative detuning to being dominated by strongly bound $|1\rangle$-$|3\rangle$ pairs at positive detuning. While the zero crossing happens at zero detuning in the non-interacting system, we find that once we have interactions, $\mathcal{S}_{z}$ crosses zero at negative detuning roughly around the point where $\Delta$ is reduced by a factor 2 compared with the BEC limit. Furthermore, increasing $\Omega/E_{F}$ does not affect the position of the zero crossing, and only changes the scale of detunings over which the crossover occurs. The behavior of the zero crossing is similar to what has been obtained for the Rabi-driven Fermi polaron [29, 31]. The in-plane magnetization $\mathcal{S}_{x}$ is also similarly peaked at the point where $\mathcal{S}_{z}=0$, but unlike the non-interacting case, it does not reach $-1$.
@@
Figure 9: Quasiparticle excitation spectrum $E_{{\bf k}\uparrow}$ (purple), $E_{{\bf k}+}$ (blue), and $E_{{\bf k}-}$ (red) given in Eqs. (26a) and (28), for equal masses $m=m_{3}$. We fix $\Omega/E_{F}=1$, $11/k_{F}a_{13}=$ 1 and $a_{23}=0$, and consider detunings that span the crossover, with $\delta/E_{F}=-3.4$ in (a), $\delta/E_{F}=-2.1$ in (b), and $\delta/E_{F}=3$ in (c). The inset in (a) shows the zoomed in area around the minimum of the quasiparticle excitation spectrum, where the dotted lines give the magnitude of the excitation gaps derived in Eq. (76). The dashed black lines are the upper non-interacting quasiparticle energies $\xi_{\bf k}+\epsilon_{+}$ with (a,c) $\mu-\mu_{3}=\epsilon_{-}$, and (b) $\mu-\mu_{3}=-2.6E_{F}$, corresponding to a lower value than $\epsilon_{-}\simeq-2.2E_{F}$.
We now turn to show additional quantities that characterize the state of the system, namely in Fig. 8 the density distributions and the pairing correlations, and in Fig. 9 the quasiparticle excitation spectrum. In both figures, we take the bare interaction strength $k_{F}a_{13}=1$, i.e., in the absence of the Rabi drive we would be on the BEC side of the bare resonance. However, according to Fig. 6, varying the detuning of the Rabi drive allows us to explore an effective BCS regime [$\delta/E_{F}=-3.4$], a crossover regime [$\delta/E_{F}=-2.1$], and a BEC regime [$\delta/E_{F}=3$].
@@
In Figs. 8(d,e,f) we instead plot the pairing correlations $\kappa_{{\bf k}i3}=\langle\hat{f}_{-{\bf k}i}\hat{f}_{{\bf k}3}\rangle$. In the BCS-like limit shown in Fig. 8(d), we can see that $\kappa_{{\bf k}13}$ and $\kappa_{{\bf k}23}$ are both substantial even though there is no direct interaction between species |$\ket{2}$ and |$\ket{3}$. For $k\lesssim k_{F}$, we have $|\kappa_{23}|>|\kappa_{13}|$ with both exhibiting peaks upon approaching the edge of the Fermi sea when $k\simeq k_{F}$, similar to the conventional Cooper pairing in the BCS limit of the two-component Fermi gas. On the other hand, for $k\gtrsim k_{F}$, $\kappa_{{\bf k}23}$ decays faster than $\kappa_{{\bf k}13}$. Here again, this behavior is expected since $\kappa_{{\bf k}13}\sim 1/k^{2}$ at large momenta [see Eq. (59)] whereas, since $\Delta_{2}=0$, the leading-order contribution to the tail of $\kappa_{{\bf k}23}$ is
@@
We also briefly comment on the other crossovers that one can obtain with a single interaction channel. First, suppose we have $k_{F}a_{13}<-1$. In this case of negative scattering length, the Rabi coupling acts to make the interactions weaker, and therefore the superfluid ground state has BCS character for all detunings. Likewise, if we start with unitarity-limited interactions, $01/k_{F}a_{13}=$ 0, we cannot induce a bound state via Rabi coupling, and as a function of detuning we have a BCS to unitary Fermi gas superfluid crossover. In Fig. 1 these cases are obtained by taking $a_{23}\to 0^{-}.$
@@ VI Conclusion
We also found the existence of an exotic excited-state solution, which manifests as a saddle point of the free energy. Notably, it exhibits superfluid transitions as well as a BCS-BEC-BCS crossover, and we discussed how these could potentially be accessed dynamically. Utilizing the free energy, we generalized the Thouless criterion and showed that both the ground and excited states possess critical temperatures that are comparable to those of the standard BCS-BEC crossover, although the excited state critical temperature is somewhat below that of the ground state. Therefore, the physics described in this work should be accessible in current experiments.
@@ Appendix A Low-density limit of the BCS ansatz
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -2\left(s_{\bf k}g_{13}\sum_{{\bf q}}c_{\bf q}\theta_{\bf q}-c_{\bf k}g_{23}\sum_{{\bf q}}s_{\bf q}\theta_{\bf q}\right)=0.$ (78b)
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\zeta_{\bf k}\left(s_{\bf k}g_{13}\sum_{{\bf q}}c_{\bf q}\theta_{\bf q}-c_{\bf k}g_{23}\sum_{{\bf q}}s_{\bf q}\theta_{\bf q}\right)=0.$ (79)
@@ B.1 Conventional two-component Fermi gas
It is insightful to remind ourselves how the Thouless criterion arises in the case of the homonuclear two-component ($\uparrow$,$\downarrow$ , ↓) Fermi gas with interaction constant $g<0$. In this case, the mean-field free-energy is given by [68]
@@
In this subsection only, we will use the notation $E_{\bf k}=\sqrt{\bar{\xi}_{\bf k}^{2}/4+\Delta^{2}}$, $\bar{\xi}_{\bf k}=\xi_{{\bf k}\uparrow}+\xi_{{\bf k}\downarrow}$ with $\xi_{{\bf k}\sigma}=\epsilon_{\bf k}-\mu_{\sigma}$. In addition, we have introduced $h=(\mu_{\uparrow}-\mu_{\downarrow})/2$ which plays the role of a Zeeman field which can split the $\uparrow$-$\downarrow$ - ↓ components. In the limit of small $\Delta$, the free energy can be expanded as
@@
$T^{\text{mb}}(0)^{-1}=\frac{1}{g}+\sum_{\bf k}\frac{1-f_{{\bf k}\uparrow}-f_{{\bf k}\downarrow}}{\bar{\xi}_{\bf k}-\mu_{\uparrow}-\mu_{\downarrow}}.$ (89)
@@ B.2 Rabi-coupled three-component case
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\Delta_{2}^{2}\underbrace{\left[\frac{1}{g_{23}}+\frac{1}{2}\sum_{\bf k}\left(\frac{\partial^{2}E_{\bf k}}{\partial\Delta_{2}^{2}}+R^{-1}\frac{\partial^{2}E_{\bf k}}{\partial\Delta_{1}\partial\Delta_{2}}\right)+\frac{1}{2}\sum_{{\bf k}j}f_{{\bf k}j}\left(\frac{\partial^{2}E_{{\bf k}j}}{\partial\Delta_{2}^{2}}+R^{-1}\frac{\partial^{2}E_{{\bf k}j}}{\partial\Delta_{1}\partial\Delta_{2}}\right)\right]_{\Delta_{1,2}=0}}_{B}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ +\underbrace{\sum_{{\bf k}}\zeta_{\bf k}\Theta(-\zeta_{\bf k})-\frac{1}{\beta}\sum_{{\bf k}j}\ln\left(1+e^{-\beta E_{{\bf k}j}|_{\Delta_{1,2}=0}}\right)}_{\mathcal{F}_{0}}+\text{higher order terms}.$ (90)
@@
$\mu_{3}-\mu+\epsilon_{-}=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \epsilon_{+}-\mu\geq 0.$ (94)
@@
The form of the quasiparticle energies above defines three distinct regions of momentum: $\epsilon_{-}\leq\bar{\xi}_{\bf k}$, $-\epsilon_{+}<\bar{\xi}_{\bf k}<-\epsilon_{-}$ and $\bar{\xi}_{\bf k}<-\epsilon_{+}$.

### v2

@@ IV.2 Thouless criterion at finite temperature
$\times\begin{cases}\exp(\frac{\pi/c^{2}}{2k_{F}a_{13}}+\frac{\pi^{2}}{m_{r}k_{F}}\frac{s^{2}}{c^{2}}I_{+}),&a_{23}=0,\\ \exp(\frac{\pi/s^{2}}{2k_{F}a_{23}}+\frac{\pi^{2}}{m_{r}k_{F}}\frac{c^{2}}{s^{2}}I_{+}),&a_{13}=0.\end{cases}$
@@ V.1.2 Magnetization
[…]
@@ V.1.3 Two- and three-body contacts
$\penalty 10000\ \penalty 10000\ +\sum_{{\bf k}{\bf p}{\bf q}}v_{\bf k}^{2}v_{{\bf p}}u_{\bf p}v_{{\bf q}}u_{\bf q}\left(c_{\bf k}^{2}s_{\bf p}s_{\bf q}+s_{\bf k}^{2}c_{\bf p}c_{\bf q}-2s_{\bf k}c_{\bf k}s_{\bf p}c_{\bf q}\right),$
@@
Figure 5: BEC-BCS crossover with Rabi-coupled fermions. (a) Effective scattering length, (b) BCS order parameter, (c) total chemical potential measured from the (Rabi-shifted) continuum. The results are plotted for parameters $1/k_{F}a_{13}=1$, $a_{23}=0$, and $\Omega/E_{F}=1,2,3$ (solid, dashed, and dotted lines). The inset in (b) shows the analytical order parameter in the BCS regime, Eq. (65).
@@
where we have used the facts that $g_{j3}\rightarrow 0^{-}$ and $g_{13}/g_{23}\rightarrow 1$ when we take $\Lambda\rightarrow\infty$ in Eq. (3). Since our ansatz only explicitly involves two-body correlations, the three-body contact is entirely determined from the two-body contacts and the densities. In this sense, the three-body contact as defined above can be considered to be “disconnected” and distinct from the thermodynamic variable introduced in Refs. [66, 67], which depends on genuine three-body correlations and on physics beyond the two-body scattering length (e.g., on the existence of a three-body parameter [68]. However, the three-body contact in Eq. (62) still has an important physical interpretation as it determines the rate of three-body losses, i.e., the loss rate is directly proportional to $\mathcal{C}^{(3)}_{123}$ (see, e.g., Ref. [69]).
We emphasize that the physics here is distinct from the case of three-component Fermi gases featuring near-resonant interactions between all components [39, 70, 71]. In that case, Efimov-like trimers are expected to be strongly bound and to play a major role in both few- and many-body physics. By contrast, the fewer resonant interactions in our work suppresses Efimov physics [72] while the strong Rabi coupling effectively reduces the system to two components. Thus we expect that the formation of Efimov trimers can be safely neglected in our scenario.
@@ V.2 BEC-BCS crossover for a single interaction channel
[…]
@@
We see that when $\delta/\Omega\gg 1$, $a_{-}$ approaches the bare scattering length $a_{13}$, and that if $_{13}>0, aa_{-}$ can exhibit a resonance (i.e., a zero crossing of $1/a_{-}$) by controlling the detuning. This behavior is illustrated in Fig. 5(a) which shows $1/k_{F}a_{-}$ as a function of detuning for a fixed bare scattering length $1/k_{F}a_{13}=1$ and different magnitudes of the Rabi drive: $\Omega/E_{F}=1$, 2, and 3.
@@
Figure 6: (a) Out-of-plane and in-plane magnetization $\mathcal{S}_{z}$ (red) and $\mathcal{S}_{x}$ (blue), respectively, and (b) disconnected three-body contact. Parameters as in Fig. 5: $1/k_{F}a_{13}=1$, $a_{23}=0$, and $\Omega/E_{F}=1,2,3$ (solid, dashed, and dotted lines).
@@
Figure 7: (a,b,c) Density distributions and (d,e,f) pairing correlations across the BEC-BCS crossover. From left to right we have $\delta/E_{F}=-3.4,-2.1,3$. The dashed black lines in (d,e,f) are obtained using the low-density approximation for $c_{\bf k}$ and $s_{\bf k}$ when $\bar{\xi}_{\bf k}\geq|\epsilon_{-}|$ [Eq. (23)] and the noninteracting coefficients [Eq. (6)] when $\bar{\xi}_{\bf k}<|\epsilon_{-}|$, substituted into Eq. (49). The results are plotted for parameters $\Omega/E_{F}=1$, $1/k_{F}a_{13}=1$, $a_{23}=0$.
We also investigate the multi-component nature of the Rabi-coupled Fermi gas, as encoded in the magnetization and three-body contact. In Fig. 6(a) we see that $\mathcal{S}_{z}$ evolves smoothly from 1-1 to +1+1 as a function of detuning. This is natural as the ground state changes from being composed of a set of weakly perturbed Fermi seas in the $|2\rangle$ and $|3\rangle$ components at large negative detuning to being dominated by strongly bound $|1\rangle$-$|3\rangle$ pairs at positive detuning. While the zero crossing happens at zero detuning in the non-interacting system, we find that once we have interactions, $\mathcal{S}_{z}$ crosses zero at negative detuning roughly around the point where $\Delta$ is reduced by a factor 2 compared with the BEC limit. Furthermore, increasing $\Omega/E_{F}$ does not affect the position of the zero crossing, and only changes the scale of detunings over which the crossover occurs. The behavior of the zero crossing is similar to what has been obtained for the Rabi-driven Fermi polaron [29, 31]. The in-plane magnetization $\mathcal{S}_{x}$ is also similarly peaked at the point where $\mathcal{S}_{z}=0$, but unlike the non-interacting case, it does not reach 1-1.
@@
Figure 8: Quasiparticle excitation spectrum $E_{{\bf k}\uparrow}$ (purple), $E_{{\bf k}+}$ (blue), and $E_{{\bf k}-}$ (red) given in Eqs. (26a) and (28), for equal masses $m=m_{3}$. We fix $\Omega/E_{F}=1$, $1/k_{F}a_{13}=1$ and $a_{23}=0$, and consider detunings that span the crossover, with $\delta/E_{F}=-3.4$ in (a), $\delta/E_{F}=-2.1$ in (b), and $\delta/E_{F}=3$ in (c). The inset in (a) shows the zoomed in area around the minimum of the quasiparticle excitation spectrum, where the dotted lines give the magnitude of the excitation gaps derived in Eq. (68). The dashed black lines are the upper non-interacting quasiparticle energies $\xi_{\bf k}+\epsilon_{+}$ with (a,c) $\mu-\mu_{3}=\epsilon_{-}$, and (b) $\mu-\mu_{3}=-2.6E_{F}$, corresponding to a lower value than $\epsilon_{-}\simeq-2.2E_{F}$.
We now turn to show additional quantities that characterize the state of the system, namely in Fig. 7 the density distributions and the pairing correlations, and in Fig. 8 the quasiparticle excitation spectrum. In both figures, we take the bare interaction strength $k_{F}a_{13}=1$, i.e., in the absence of the Rabi drive we would be on the BEC side of the bare resonance. However, according to Fig. 5, varying the detuning of the Rabi drive allows us to explore an effective BCS regime $\delta/E_{F}=-3.4$], a crossover regime $\delta/E_{F}=-2.1$], and a BEC regime $\delta/E_{F}=3$].
@@
[…]
@@
We also briefly comment on the other crossovers that one can obtain with a single interaction channel. First, suppose we have $k_{F}a_{13}<-1$. In this case of negative scattering length, the Rabi coupling acts to make the interactions weaker, and therefore the superfluid ground state has BCS character for all detunings. Likewise, if we start with unitarity-limited interactions, $1/k_{F}a_{13}=0$, we cannot induce a bound state via Rabi coupling, and as a function of detuning we have a BCS to unitary Fermi gas superfluid crossover. In Fig. 1 these cases are obtained by taking $a_{23}\to 0^{-}.$
@@ VI Conclusion
We also found the existence of an exotic excited-state solution, which manifests as a saddle point of the free energy. Notably, it exhibits superfluid transitions as well as a BCS-BEC-BCS crossover, and we discussed how these could potentially be accessed dynamically. Utilizing the free energy, we generalized the Thouless criterion and showed that both the ground and excited states possess critical temperatures that are comparable to those of the standard BCS-BEC crossover, although the excited state critical temperature is somewhat below that of the ground state. Therefore, the physics described in this work should be accessible in current experiments. To distinguish between the ground and excited branches, two key observables can be used. First, the magnetization of the system, which depends on the detuning, can be reconstructed using spin-resolved imaging, offering a clear signature of the underlying state. Second, measuring the pairing gap across the superfluid transition—following the methodology demonstrated in Ref. [77]—would provide a direct probe of the many-body state and its evolution with temperature.
@@ Acknowledgements.
VII Data Availability
The data that support the findings of this article are openly available [90].
@@ Appendix A Low-density limit of the BCS ansatz
$\penalty 10000\ \penalty 10000\ \penalty 10000\ -2\left(s_{\bf k}g_{13}\sum_{{\bf q}}c_{\bf q}\theta_{\bf q}-c_{\bf k}g_{23}\sum_{{\bf q}}s_{\bf q}\theta_{\bf q}\right)=0.$ (70b)
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ +\zeta_{\bf k}\left(s_{\bf k}g_{13}\sum_{{\bf q}}c_{\bf q}\theta_{\bf q}-c_{\bf k}g_{23}\sum_{{\bf q}}s_{\bf q}\theta_{\bf q}\right)=0.$ (71)
@@ B.1 Conventional two-component Fermi gas
It is insightful to remind ourselves how the Thouless criterion arises in the case of the homonuclear two-component ($\uparrow,\downarrow$) Fermi gas with interaction constant $g<0$. In this case, the mean-field free-energy is given by [73]
@@
In this subsection only, we will use the notation $E_{\bf k}=\sqrt{\bar{\xi}_{\bf k}^{2}/4+\Delta^{2}}$, $\bar{\xi}_{\bf k}=\xi_{{\bf k}\uparrow}+\xi_{{\bf k}\downarrow}$ with $\xi_{{\bf k}\sigma}=\epsilon_{\bf k}-\mu_{\sigma}$. In addition, we have introduced $h=(\mu_{\uparrow}-\mu_{\downarrow})/2$ which plays the role of a Zeeman field which can split the $\uparrow$-$\downarrow$ components. In the limit of small $\Delta$, the free energy can be expanded as
@@
$T^{\text{mb}}(0)^{-1}=\frac{1}{g}+\sum_{\bf k}\frac{1-f_{{\bf k}\uparrow}-f_{{\bf k}\downarrow}}{\bar{\xi}_{\bf k}}.$ (81)
@@ B.2 Rabi-coupled three-component case
$\penalty 10000\ \penalty 10000\ \penalty 10000\ -\Delta_{2}^{2}\underbrace{\left[\frac{1}{g_{23}}+\frac{1}{2}\sum_{\bf k}\left(\frac{\partial^{2}E_{\bf k}}{\partial\Delta_{2}^{2}}+R^{-1}\frac{\partial^{2}E_{\bf k}}{\partial\Delta_{1}\partial\Delta_{2}}\right)+\frac{1}{2}\sum_{{\bf k}j}f_{{\bf k}j}\left(\frac{\partial^{2}E_{{\bf k}j}}{\partial\Delta_{2}^{2}}+R^{-1}\frac{\partial^{2}E_{{\bf k}j}}{\partial\Delta_{1}\partial\Delta_{2}}\right)\right]_{\Delta_{1,2}=0}}_{B}$
$\penalty 10000\ \penalty 10000\ +\underbrace{\sum_{{\bf k}}\zeta_{\bf k}\Theta(-\zeta_{\bf k})-\frac{1}{\beta}\sum_{{\bf k}j}\ln\left(1+e^{-\beta E_{{\bf k}j}|_{\Delta_{1,2}=0}}\right)}_{\mathcal{F}_{0}}+\text{higher order terms}.$ (82)
@@
$\mu_{3}-\mu+\epsilon_{-}=0\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \penalty 10000\ \epsilon_{+}-\mu\geq 0.$ (86)
@@
The form of the quasiparticle energies above defines three distinct regions of momentum: $\epsilon_{-}\leq\bar{\xi}_{\bf k}$, $\epsilon_{+}<\bar{\xi}_{\bf k}<-\epsilon_{-}$ and $\bar{\xi}_{\bf k}<-\epsilon_{+}$.